OpenStudy (anonymous):

ABCD is a rectangle. find the length of each diagonal. AC=2(5a+1) BD=2(a+1)

OpenStudy (anonymous):

I got 0

OpenStudy (anonymous):

Impossiple for a real rectangle

OpenStudy (anonymous):



OpenStudy (anonymous):

I assume it looks like that

OpenStudy (anonymous):

we know that

OpenStudy (anonymous):

2(5s+1)=2(a+1)

OpenStudy (anonymous):

I also think you might have a typo 2(5s+1)=2(a+1) ^ ^ I think either both s or both a

OpenStudy (anonymous):

both a

OpenStudy (anonymous):

Either way if both s and a are the same you get 0 Put you plug it back in to either side of the equation 2(0+1) 0+1=1 I get 1

OpenStudy (anonymous):

You missunderstand G. a=0 :P but not the diagonals the diagonals = 2(5a+1) or 2(a+1) plug in 0 at and you get 2(0+1) 0+1=1 or 2(5*0+1) 0+1=1

OpenStudy (anonymous):

YOU MADE ME GET THE WRONG ANSWER

OpenStudy (anonymous):

what is right answer then?

OpenStudy (anonymous):

who knws im not on this problem any more hint its closed

OpenStudy (anonymous):

2(5*0+1) 0+2 2 i distributed wrong :( It was 2 sorry

OpenStudy (anonymous):

OMG IM NOT ON THIS PROBLEM ANYMORE